Two straight, parallel, current-carrying wires separated by a distance of 5cm contain the same 1000 A current, flowing in similar directions. Determine the force per unit length exerted by the two wires.



1
Expert's answer
2021-05-05T16:32:46-0400


Given:

2 straight current carrying wire

i=1000 A (in both)

direction= same ( therefore attractive force)

distance b/w them (d) = 5 cm


for calculating we use EMF formula

as, magnetic force will be generated b/w both the wires.


"\boxed {F={\mu_o( i_1 )(i_2 \over2\pi d}}"


but, here i1=i2

therefore,


"\boxed {F={\mu_o i^2 \over2\pi d}}"


after putting values we get,


"\boxed {F={(4 \pi \times 10^{-7})(1000)^2\over 2\pi (5 \times10^{-2})}}"


"\boxed {F=4 Nm^{-1}}"